OpenStudy (anonymous):

Factor 13c^2-52. Can someone help me complete this problem?

OpenStudy (anonymous):

13c^2 -52 = 13 ( c^2 -4) = 13 ( c+2) ( c-2)

OpenStudy (anonymous):

What method did you use to solve this?

OpenStudy (anonymous):

use this formula: a^2 - b^2 = ( a+b) ( a-b)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem :)